translate to a system of equations and solve.
the sum of two numbers is 30. their difference is 40. find the numbers.

Answer by xcentaur(357)

let the numbers be x and y


x+y=30
x-y=40


adding the two eqns we get
2x=70
x=35


y=30-x
y=30-35
y=-5


Hence the numbers are: 35,-5